# Efficient Neuro-Symbolic Learning of Constraints & Objective

### v1

@@ 5.1. Learning how to play logic puzzles
The NP-complete Sudoku problem is a classical logical reasoning problem that has been repeatedly used as a benchmark in a “learning to reason” context (Palm et al., 2018; Amos and Kolter, 2017; Wang et al., 2019; Brouard et al., 2020; Defresne et al., 2023; Yang et al., 2023; Li et al., 2023b; Bessiere et al., 2023; Ye et al., 2025). The task is to learn how to solve new Sudoku grids from a set of solved grids, without knowing the game rules. On this task and two variants, we show that the E-PLL enables to learn the exact rules of the puzzles with high time-and-data-efficiency.
@@ 5.1.1. Experimental setting
Given samples $(\omega^{\ell},\mathbf{y}^{\ell})$ of initial and solved Sudoku grids, we want to learn how to solve new grids. Sudoku players know that Sudoku grids can be more or less challenging. As one could expect, it is also harder to train how to solve hard grids than easy grids (Brouard et al., 2020). We use the number of initially filled cells (hints) as a proxy for the problem’s hardness, a grid with few hints being hard. The minimal number of hints required to define a single solution is 17, defining the hardest single-solution Sudoku grids. We use the RRN data set (Palm et al., 2018), composed of single-solution grids with $17$ to $34$ hints. We use $100$ grids for training, and $32$ for validation (all hardness). As in (Palm et al., 2018), we test on the hardest 17-hints instances, $1,000$ in total.
@@
Published Sudoku grids have only one solution. We also consider a Sudoku benchmark used in (Nandwani et al., 2021), where each grid has more than one solution. For each grid, the set of solutions is only partially accessible during training: at most $5$ of them are present in the training set. The aim is to be able to predict any of the feasible solutions (all of them are known for testing). We use $100$, $32$, and $256$ grids of the data set from (Nandwani et al., 2021) respectively for training, validating, and testing. During training, one of the $5$ available solutions for a given grid is randomly selected.
@@
A $9\times 9$ Sudoku grid is represented as $81$ cell coordinates with a possible hint when available. Each cell is represented by a GM variable with domain $\{1,\ldots,9\}$. For $N$, we reuse the same architecture (Defresne et al., 2023), but we drastically reduce the number of parameters. We use a Multi-Layer Perceptron (MLP) with $4$ hidden layers of $64$ neurons and residual connections (He et al., 2016) every $2$ layers. It receives the pairs of coordinates of pairs of cells $(Y_{i},Y_{j})$ and predicts all pairwise cost matrices $N(\omega)[i,j]$. Hints are used to set the values of their corresponding variable in $N(\omega)$. Performances are measured by the percentage of correctly filled grids; no partial credit is given for individual digits.
@@
is another grid-based puzzles were the solution must respect the same rules as Sudoku plus inequalities between some pairs of cells (see Figure 3). We generated our own dataset, composed of $5\times 5$ grids (for details, see annex C.1). We used $1000$ training grids, $64$ for validation and $200$ for testing. The grid representation is similar to Sudoku: each pair of cells is represented by its coordinates (row and column indices) and an additional feature representing the inequality ($1$ if $i>j$, 1-1 is $i<j$ and 0 if no inequality). The same neural net as for Sudoku is used.
@@ 5.1.2. Test results
Since the E-PLL never compares a solver-produced solution to the provided solution $\mathbf{y}$, it is not sensitive to the existence of many solutions. Therefore, training on the dataset with multiple solutions yields similar results: one of the feasible solution is predicted for 100% of the test grids. We come to a similar observation with Futoshiki. For inference, a threshold is applied on all the cost functions: all costs below $1$ are set to 0. All of the test grids were correctly solved.
In Table 2, we compare our results with related approaches that learn how to solve Sudoku. Pure Deep Learning methods, Recurrent Relational Network (RRN) (Palm et al., 2018), Recurrent Transformer (Yang et al., 2023), and Denoising Diffusion Probabilistic Models (DDPMs) (Ye et al., 2025), require orders of magnitude more data and fail to solve some of the hardest puzzles. The accuracy of Recurrent Transformers drops by roughly 20% when the amount of data is reduced to 9,000 samples. While DDPMs (Ye et al., 2025) solve Sudokus (from an easy Kaggle dataset) quite reliably, they completely fail on hard Sudokus. Adding a convex relaxation of Max-SAT reasoning optimization layer, SATNet (Wang et al., 2019) becomes much more data-efficient. Still, it fails to solve some of the easy grids, and its accuracy drops significantly on hard grids, below that of pure DL approaches.
Using a discrete solver alone (Bessiere et al., 2023), or combined with ML (Brouard et al., 2020), it is possible to solve all test instances reliably from properly learned rules, offering far more efficient training than DL-based approaches. However, these non-end-to-end differentiable approaches cannot directly exploit natural inputs, as in the Visual Sudoku or protein design tasks described below. Hybrid approaches combine both DL and exact reasoning. A follow-up work (Li et al., 2023a) of SATNet, extracts explicit logical rules from SATNet, enabling reliable solving on $4\times 4$ grids. On $9\times 9$ grids, it learns hundreds of thousands of clauses, leading to unsolvable instances. Approaches that embed an exact solver during training, as the structured Perceptron/Hinge losses, can solve any grid (Defresne et al., 2023), at the cost of an excruciatingly long training time. Overall, our approach stands out for its perfect reliability, its data-efficiency, and its low training time.
@@
DL RRN (Palm et al., 2018) 96.6% 17 180,000 200k >50>50
Rec. Trans. (Yang et al., 2023) 96.796.7% 17 180,000 211k >50>50
@@
DDPM 0.2% 17 - - -
Relax+DL SATNet (Wang et al., 2019) 95.1–99.8% 36.2 9,000 600k 2.9
@@
CO (Bessiere et al., 2023) 100% - 200 - 0.01
CO + ML (Brouard et al., 2020) 100% 17 9,000 - 1.5
CO+DL Hinge (Defresne et al., 2023) 100% 17 1,000 180k >>50
@@
Table 2. Accuracies of related works. They are sorted by type of approach: pure Deep Learning (DL), reasoning (CO for combinatorial optimization), and relaxation of reasoning (Relax), which can be combined with ML or DL. The ’# hints’ gives the average hardness of the test set. Param. is the number of parameters of the neural network.
@@ 5.1.3. Retrieving exact constraints
After learning, logical constraints can be retrieved with two alternative strategies. When constraints only need to be learned, a threshold can be applied (as done for Futoshiki): learnt costs below the threshold are set to 0 and costs above are set to $\infty$ (hard constraint). The second strategy is cost function hardening (Brouard et al., 2020), which has the advantage of requiring no parameter and also of preserving the learned objective (if needed). Non-zero learned costs are considered in decreasing order and set to $\infty$ if the corresponding value combination is not observed in any of the training set. This is repeated until a contradiction is found.
@@ 5.2. Visual Sudoku
Our data set is obtained from the symbolic Sudoku data set by replacing hints with arbitrarily selected MNIST images, as in (Brouard et al., 2020; Topan et al., 2021). We use grids from the RRN dataset ($9,000$ for training, $64$ for validation), as they are much more challenging than the SATNet dataset (average $36.2$ hints) (Wang et al., 2019). The test set contains 100 grids of each difficulty (from $17$ to $34$ initial hints).
@@
The original Visual Sudoku task (Wang et al., 2019) presented a subtle form of data leakage (Chang et al., 2020): each hint digit image in the input $\omega$ had a corresponding label in the output $\mathbf{y}$, and it was possible to ground every image into its specific meaning directly. In the ungrounded Visual Sudoku task, the training set contains no information on the labels of the cells with hints’ images. We handle this as missing data, using imputation to predict complete assignments from incomplete assignments $\mathbf{y}$ in the training set. During the first epochs, the predicted rules are mostly random, making the imputation longer, Therefore, we first restrict training to grids with few initial hints: only $17$ on first epoch, then less than $20$ on the second, $30$ on the third and so on until all training grids are included. Moreover, we decay the learning rate of both neural nets by dividing it by $10$ at epochs $6$ and $8$, training for $20$ epochs.
@@
After visual Sudoku training, we assessed the accuracy of the CNN alone on MNIST digits, the percentage of correctly-filled grids when no correction of MNIST classification is allowed, and the final percentage of correct grids with corrections by the solver. We only compare methods tackling the grounded problem, where no label is available for hints. The resulting perception architecture is not as accurate as competitors, with a 1% lower accuracy, leading to far fewer properly filled grids. Yet, the Sudoku rules are learnt properly, enabling the solver to correct 2020% of the grids eventually reaching a competitive $93.3\%$ of visual grids solved, with a training time reduced by 3232%.
@@ 5.3. Application in a DFL setting with Min-Cut and Max-Cut
Although our method was originally designed to learn on SOP datasets $(\omega^{\ell},\mathbf{y}^{\ell})$, with no intermediate supervision, it can also be applied to DFL tasks with known constraints. We designed Min-Cut/Max-Cut DFL tasks where a graph $G=(V,E)$ is given, where each edge $e\in E$ is associated with an image of a stone, wood, or rope bridge, with decreasing respective capacities of 5, 2, and 1.
@@
To build training samples, we use the graph of a truncated icosahedron, an Archimedean solid with a planar graph, and select a source vertex randomly. The sink node is set to the opposite vertex in the solid. To avoid trivial cuts, edges incident to the source and sink nodes are associated with stone-bridge images. The remaining edges are assigned a random image/capacity (stone, wood, or rope).
The Min/Max cut problems can be easily encoded in a GM where a Boolean variable is associated with every vertex, indicating its side of the cut. For every edge $e$ in $E$, a cost function proportional to a $2\times 2$ identity matrix $I_{2}$ (Max-Cut) or $(1-I_{2})$ matrix (Min-Cut) is used. The multiplicative factor represents the cost to pay, if the edge is cut (Min-Cut) or not (Max-Cut). The DFL dataset $(\omega^{\ell},\mathbf{c}^{\ell})$ is turned into a SOP dataset $(\omega^{\ell},y^{\ell})$ by solving each instance $(\omega^{\ell},\mathbf{c}^{\ell})$ with the exact GM solver toulbar2 (Hurley et al., 2016). The $(1-I_{2})$ matrix being submodular, Min-Cut is solved efficiently (Cooper et al., 2010). Max-Cut is NP-hard.
@@
We compared the E-PLL with the pioneer DFL SPO+ loss (Elmachtoub and Grigas, 2022), trained with a learning rate of $10^{-4}$. As shown in Figure 5, the E-PLL and SPO+ minimize regret until it reaches 0 for both tasks. SPO+ requires the true solution $\mathbf{y}^{\ell}$ for each instance, so it does not save the time needed to convert the training set from a DFL to an SOP setting. Instead, it requires one extra optimization using $2\hat{\mathbf{c}}-\mathbf{c}$ as the objective function, for every sample during training. We observe that the E-PLL converges faster and almost monotonically. This is quite remarkable given that 1) it uses a weaker training signal ($y^{\ell}$) than SPO+, which exploits both $c^{\ell}$ and $y^{\ell}$, and 2) that it also does not explicitly minimize regret. Deciding which approach to use in practice also depends on the problem : some problems may be easier to model and may be more efficiently solved using (I)LP – as done with SPO+ – or using pairwise GMs (Hurley et al., 2016).
Figure 5. Test regret on MinCut (left) and MaxCut (right) tasks, averaged over 10 runs (different initializations).
@@ 5.4.1. Problem definition
The problem of designing proteins has similarities with solving Sudokus (Strokach et al., 2020). Proteins are linear macro-molecules defined by a sequence of smaller proteins called amino acids. Proteins usually fold into a specific 3D structure which enables specific biological or biochemical functions. Designed proteins have applications in health and green chemistry, among others (Kuhlman and Bradley, 2019). To design new proteins, with new functions or enhanced properties, one usually starts from an input backbone structure, matching the target functions, and predicts a sequence of amino acids that will fold onto the target structure.
Considering an input protein structure as a Sudoku grid, each amino acid corresponds to a cell and must be chosen among the $20$ natural amino acids instead of $9$ digits. The structure is predominantly determined by inter-atomic forces, which drive folding into a minimum energy geometry: the most usual approach of the protein design problem is as an NP-hard unconstrained energy minimization problem (Pierce and Winfree, 2002; Allouche et al., 2014). Inter-atomic forces are influenced by relative distances and atomic natures, which implies that the final interactions inside a protein depend on the geometry of the input structure: we will therefore use this geometry in the natural inputs $\omega$, as we did with the (fixed) Sudoku grid geometry.
Compared to Sudoku, protein design raises several challenges: 1) instances have variable geometries, which is tackled by a dedicated neural network. 2) They may contain several hundred amino acids (i.e, variables) or more, requiring an approach able to scale to large instances. 3) No ground-truth cost function is available. In fact, proteins are subjected to many-body interactions that cannot be directly represented in a pairwise GM. We aim to learn a better approximation than the score functions used for design, such as in Rosetta (Park et al., 2016). 4) Many sequences can fit the same input backbone structure, similarly to the one-of-many solutions setting. The parallel between Sudoku and protein design is illustrated in Figure 2.

### v3

@@ 5.1. Learning how to Play Logic Puzzles
The NP-complete Sudoku problem is a classical logical reasoning problem that has been repeatedly used as a benchmark in a “learning to reason” context (RRN, 54, 5, 70, 13, 21, 73, 36, 10, 75). The task is to learn how to solve new Sudoku grids from a set of solved grids, without knowing the game rules. On this task and two variants, we show that the E-PLL enables to learn the exact rules of the puzzles with high time-and-data-efficiency.
@@ 5.1.1. Experimental Setting
Given samples $(\omega^{\ell},\mathbf{y}^{\ell})$ of initial and solved Sudoku grids, we want to learn how to solve new grids. Sudoku players know that Sudoku grids can be more or less challenging. As one could expect, it is also harder to train how to solve hard grids than easy grids (brouard2020pushing, 13). We use the number of initially filled cells (hints) as a proxy for the problem’s hardness, a grid with few hints being hard. The minimal number of hints required to define a single solution is 17, defining the hardest single-solution Sudoku grids. We use the RRN data set (RRN, 54), composed of single-solution grids with $17$ to $34$ hints. We use $100$ grids for training, and $32$ for validation (all hardness). As in (RRN, 54), we test on the hardest 17-hints instances, $1,000$ in total.
Recently, a new Sudoku-Extreme dataset (wang2025hierarchical, 69) was assembled based on the number of backtracks needed by the tdoku solver (github.com/t-dillon/tdoku) to solve instances. In (wang2025hierarchical, 69), $1,000$ extreme grids are used and augmented $1,000$ times by a dedicated augmentation procedure, reaching a $1,001,000$ grids training set. We use $10$ extreme grids from the same data set, augmented $99$ times only, with same augmentation code, to reach a final training set of $1,000$ grids for training. For validation, we use the same validation subset as (wang2025hierarchical, 69).
@@
Published Sudoku grids have only one solution. We also consider a Sudoku benchmark used in (1ofMany, 51), where each grid has more than one solution. For each grid, the set of solutions is only partially accessible during training: at most $5$ of them are present in the training set. The aim is to be able to predict any of the feasible solutions (all of them are known for testing). We use $100$, $32$, and $256$ grids of the data set from (1ofMany, 51) respectively for training, validating, and testing. During training, one of the $5$ available solutions for a given grid is randomly selected.
@@
A $9\times 9$ Sudoku grid is represented as $81$ cell coordinates with a possible hint when available. Each cell is represented by a GM variable with domain $\{1,\ldots,9\}$. For $N$, we reuse the same architecture (epll, 21), but we drastically reduce the number of parameters. We use a Multi-Layer Perceptron (MLP) with $4$ hidden layers of $64$ neurons and residual connections (He_2016_CVPR, 26) every $2$ layers. It receives the pairs of coordinates of pairs of cells $(Y_{i},Y_{j})$ and predicts all pairwise cost matrices $N(\omega)[i,j]$. Hints are used to set the values of their corresponding variable in $N(\omega)$. Performances are measured by the percentage of correctly filled grids; no partial credit is given for individual digits.
@@
is another grid-based puzzle. As for Sudoku, no digit can be repeated twice in a row or column. In addition, the solution must respect inequalities between some pairs of cells (see Figure 3). We generated our own dataset, composed of $5\times 5$ grids (for details, see annex C.1). We used $1,000$ training grids, $64$ for validation and $200$ for testing. The grid representation is similar to Sudoku: each pair of cells is represented by its coordinates (row and column indices) and an additional feature representing the inequality ($1$ if $i>j$, 1-1 is $i<j$ and 0 if no inequality). The same neural net as for Sudoku is used.
@@ 5.1.2. Test Results
Since the E-PLL never compares a solver-produced solution to the provided solution $\mathbf{y}$, it is not sensitive to the existence of many solutions. Therefore, training on the dataset with multiple solutions yields similar results: one of the feasible solutions is predicted for 100% of the test grids. We come to a similar observation with Futoshiki. For inference, a threshold is applied on all the cost functions: all costs below $1$ are set to 0. All of the test grids were correctly solved.
In Table 2, we compare our results with related approaches that learn how to solve Sudoku. Pure Deep Learning methods, Recurrent Relational Network (RRN) (RRN, 54), Recurrent Transformer (Yang2023learning, 73), a Hierarchical variant (wang2025hierarchical, 69) and Denoising Diffusion Probabilistic Models (DDPMs) (ye2025beyond, 75), require orders of magnitude more data and fail to solve some of the hardest puzzles. The accuracy of Recurrent Transformers drops by roughly 20% when the amount of data is reduced to 9,000 samples. While DDPMs (ye2025beyond, 75) solve Sudokus (from an easy Kaggle dataset) quite reliably, they completely fail on hard Sudokus. Adding a convex relaxation of Max-SAT reasoning optimisation layer, SATNet (satnet, 70) becomes much more data-efficient. Still, it fails to solve some of the easy grids, and its accuracy drops significantly on hard grids, below that of pure DL approaches.
Using a discrete solver alone (ijcai2023p208, 10), or combined with ML (brouard2020pushing, 13), it is possible to solve all test instances reliably from properly learned rules, offering far more efficient training than DL-based approaches. This also extends to the Sudoku-extreme dataset, where only $10$ grids with $99\times$ augmentation suffice to reach a 100% accuracy, compared to $1,000$ grids with $1,000\times$ augmentation needed to reach 55% accuracy by the Hierachical Reasoning Model (wang2025hierarchical, 69).
However, these non-end-to-end differentiable approaches cannot directly exploit natural inputs, as in the Visual Sudoku or protein design tasks described below. Hybrid approaches combine both DL and exact reasoning. A follow-up work (li2023learning, 37) of SATNet, extracts explicit logical rules from SATNet, enabling reliable solving on $4\times 4$ grids. On $9\times 9$ grids, it learns hundreds of thousands of clauses, leading to unsolvable instances. Approaches that embed an exact solver during training, as the structured Perceptron/Hinge losses, can solve any grid (epll, 21), at the cost of an excruciatingly long training time. Overall, our approach stands out for its perfect reliability, its data-efficiency, and its low training time.
Table 2. Accuracies of related works. They are sorted by type of approach: pure Deep Learning (DL), reasoning (CO for combinatorial optimisation), and relaxation of reasoning (Relax), which can be combined with ML or DL. The ’# hints’ gives the average hardness of the test set. Param. is the number of parameters of the neural network. HRM train-time is extracted from (wang2025hierarchical, 69) running on a fast RTX 4070M as their code requires a recent GPU.
@@
DL RRN (RRN, 54) $96.6\%$ 17 180,000 200k >50>50
Rec. Trans. (Yang2023learning, 73) 96.796.7% 17 180,000 211k >50>50
@@
DDPM $0.2\%$ 17 - - -
HRM (wang2025hierarchical, 69) $55\%$ 24.8 $1,000\times 1,001$ 27M ¿10h
@@
[…]
CO (ijcai2023p208, 10) 100% - 200 - 0.01
CO + ML (brouard2020pushing, 13) 100% 17 9,000 - 1.5
CO+DL Hinge (epll, 21) 100% 17 1,000 180k >>50
@@
E-PLL (HRM dataset) 100% 24.8 $\textbf{10}\times 100$ 22k 0.04
@@ 5.1.3. Retrieving Exact Constraints
After learning, logical constraints can be retrieved with two alternative strategies. When constraints only need to be learned, a threshold can be applied (as done for Futoshiki): learnt costs below the threshold are set to 0 and costs above are set to $\infty$ (hard constraint). The second strategy is cost function hardening (brouard2020pushing, 13), which has the advantage of requiring no parameter and also of preserving the learned objective (if needed). Non-zero learned costs are considered in decreasing order and set to $\infty$ if the corresponding value combination is not observed in any of the training set. This is repeated until a contradiction is found.
@@ 5.2. Visual Sudoku
Our data set is obtained from the symbolic Sudoku data set by replacing hints with arbitrarily selected MNIST images, as in (brouard2020pushing, 13, 67). We use grids from the RRN dataset ($9,000$ for training, $64$ for validation), as they are much more challenging than the SATNet dataset (average $36.2$ hints) (satnet, 70). The test set contains 100 grids of each difficulty (from $17$ to $34$ initial hints).
@@
The original Visual Sudoku task (satnet, 70) presented a subtle form of data leakage (assessingSATNet, 15): each hint digit image in the input $\omega$ had a corresponding label in the output $\mathbf{y}$, and it was possible to ground every image into its specific meaning directly. In the ungrounded Visual Sudoku task, the training set contains no information on the labels of the cells with hints’ images. We handle this as missing data, using imputation to predict complete assignments from incomplete assignments $\mathbf{y}$ in the training set. During the first epochs, the predicted rules are mostly random, making the imputation longer. Therefore, we first restrict training to grids with few initial hints: only $17$ on first epoch, then less than $20$ on the second, $30$ on the third and so on until all training grids are included. Moreover, we decay the learning rate of both neural nets by dividing it by $10$ at epochs $6$ and $8$, training for $20$ epochs.
@@
After visual Sudoku training, we assessed the accuracy of the CNN alone on MNIST digits, the percentage of correctly-filled grids when no correction of MNIST classification is allowed, and the final percentage of correct grids with corrections by the solver. In Table 3, we only compare methods tackling the grounded problem, where no label is available for hints. The resulting perception architecture is not as accurate as competitors, with a 1% lower accuracy, leading to far fewer properly filled grids. Yet, the Sudoku rules are learnt properly, enabling the solver to correct digit missclassification inconsistent with the rules. In fine, 2020% of the grids are corrected, and the E-PLL reaches a competitive $93.3\%$ of visual grids solved, with a training time reduced by 3232%.
@@ 5.3. Application in a DFL Setting with Min-Cut and Max-Cut
Although our method was originally designed to learn on SOP datasets $(\omega^{\ell},\mathbf{y}^{\ell})$, with no intermediate supervision, it can also be applied to DFL tasks with known constraints. To illustrate the versatility of our approach, we designed a new DFL task based on the Min-Cut and Max-Cut problems. A graph $G=(V,E)$ is given, where each edge $e\in E$ is associated with a capacity. To include natural inputs, each capacity is represented by an image of a bridge, either stone, wood or rope,, with decreasing respective capacities of 5, 2, and 1. The goal is to find a cut of the graph with minimum or maximum capacity.
@@
To build training samples, we use the graph of a truncated icosahedron, an Archimedean solid with a planar graph, and select a source vertex randomly. The sink node is set to the opposite vertex in the solid. To avoid trivial cuts, edges incident to the source and sink nodes are associated with stone-bridge images. The remaining edges are assigned a random image/capacity (stone, wood, or rope). All images are noised before normalization with a Gaussian noise of mean 0 and standard deviation of $10$ (for 256 levels, 8 bit images).
The Min/Max cut problems can be easily encoded in a GM where a Boolean variable is associated with every vertex, indicating its side of the cut. For every edge $e$ in $E$, a cost function proportional to a $2\times 2$ identity matrix $I_{2}$ (Max-Cut) or $(1-I_{2})$ matrix (Min-Cut) is used. The multiplicative factor represents the cost to pay, if the edge is cut (Min-Cut) or not (Max-Cut). The DFL dataset $(\omega^{\ell},\mathbf{c}^{\ell})$ is turned into a SOP dataset $(\omega^{\ell},y^{\ell})$ by solving each instance $(\omega^{\ell},\mathbf{c}^{\ell})$ with the exact GM solver toulbar2 (hurley2016multi, 29). The $(1-I_{2})$ matrix being submodular, Min-Cut, the dual of Max-flow, is solved efficiently (cooper2010soft, 17). Max-Cut is NP-hard.
@@
We compared the E-PLL with the pioneer DFL SPO+ loss (SPO, 23), both trained with a learning rate of $10^{-3}$. SPO+ requires the true solution $\mathbf{y}^{\ell}$ for each instance, so it does not save the time needed to convert the training set from a DFL to an SOP setting. Instead, it requires one extra optimisation using $2\hat{\mathbf{c}}-\mathbf{c}$ as the objective function, for every sample during training. On larger graphs, SPO+ training time may become excruciatingly slow.
As shown in Figure 5, both the E-PLL and SPO+ minimize test regret over training iterations. On the Max-Cut task, the E-PLL converges in fewer iterations than SPO+, but it needs more for Min-Cut. This is quite remarkable given that 1) the E-PLL uses a weaker training signal ($y^{\ell}$) than SPO+, which exploits both $c^{\ell}$ and $y^{\ell}$, and 2) that it also does not explicitly minimize regret. Deciding which approach to use in practice also depends on the problem : some problems may be easier to model and may be more efficiently solved using (I)LP – as done with SPO+ – or using pairwise GMs (hurley2016multi, 29).
Figure 5. Test regret vs number of iterations on Min-Cut (left) and Max-Cut (right) tasks, averaged over 10 runs (different initializations).
@@ 5.4.1. Problem Definition
The problem of designing proteins has similarities with solving Sudokus (STROKACH2020ProteinSolver, 66). Proteins are linear macro-molecules defined by a sequence of smaller proteins called amino acids. Proteins usually fold into a specific 3D structure which enables specific biological or biochemical functions. Designed proteins have applications in health and green chemistry, among others (kuhlman2019advances, 34). To design new proteins, with new functions or enhanced properties, one usually starts from an input backbone structure, matching the target functions, and predicts a sequence of amino acids that will fold onto the target structure.
Considering an input protein structure as a Sudoku grid, each amino acid corresponds to a cell and must be chosen among the $20$ natural amino acids instead of $9$ digits. The structure is predominantly determined by inter-atomic forces, which drive folding into a minimum energy geometry: the most usual approach of the protein design problem is as an NP-hard unconstrained energy minimization problem (pierce2002protein, 56, 4). This problem is defined using scoring functions, such as in the widely-used Rosetta (RosettaDecoy, 55), to approximate the laws of physics describing the interactions between amino acids. Inter-atomic forces are influenced by relative distances and atomic natures, which implies that the final interactions inside a protein depend on the geometry of the input structure: we will therefore use this geometry in the natural inputs $\omega$, as we did with the (fixed) Sudoku grid geometry.
Compared to Sudoku, protein design raises several challenges: 1) instances have variable geometries, which is tackled by a dedicated neural network. 2) They may contain several hundred amino acids (i.e, variables) or more, requiring an approach able to scale to large instances. 3) No ground-truth cost function is available. In fact, proteins are subjected to many-body interactions that cannot be directly represented in a pairwise GM. We aim to learn a better approximation than the score functions used for design, such as in Rosetta . 4) Many sequences can fit the same input backbone structure, similarly to the one-of-many solutions setting. The parallel between Sudoku and protein design is illustrated in Figure 2.